1) In twenty-two years, Alexis' age will be two hundred ninety-seven less than four times the age of Alyssa.
2) Alexis' age is one hundred sixty-two less than three times the age of Alyssa.
What are their ages?

Answer by DeepaEvangeline(15)

Let age of Alexis be x
Let age of Alyssa be y
According to 1)
x+22 = 4(y+22) - 297
x-4y = -231 --- (1)
According to 2)
x = 3y - 162
x-3y = -162 --- (2)
eqn(2) - eqn(1) gives
y = 69
Substituting the value of y = 69 in eqn(2) gives
x-207 = -162 => x = 207-162 = 45
Alexis age is 45
Alyssa age is 69